OpenStudy (anonymous):

Let a,b,c\in R, If for all \epsilon\in(0,1), a +c\products \epsilon\leq b, then a\leq b. How to prove?

OpenStudy (anonymous):

Suppose a > b, then a -b >0. Suppose that c >0 Let \[ \epsilon =\inf \{ \frac 1 2, \frac {a-b}{5 c}t\}\\ \epsilon \in (0,1)\\ \epsilon <\frac {a-b}c \\ c\,\epsilon <a -b\\ a +c\,\epsilon> a -c\,\epsilon >b\ \] a contradiction. Try to do it with c<0

OpenStudy (kinggeorge):

For clarity, is it Let \(a,b,c \in \mathbb{R}\). If for all \(\epsilon\in(0,1), \;\; a +c\cdot \epsilon\leq b\), then \(a\leq b\). Or Let \(a,b,c \in \mathbb{R}\). If for all \(\epsilon\in(0,1), \;\; (a +c)\cdot \epsilon\leq b\), then \(a\leq b\).

OpenStudy (anonymous):

the t should not be in the inf in my post above.